# Kiet A. NguyenAdheesh JuvekarTianjiao YuMuntasir WahedIsmini Lourentzou

### v1

@@ 3.3 Correspondence Extraction Module (CEM)
Figure 3: Overview of our Correspondence Extraction Module and Adaptation Modules. In Calico, $k$ CAMs are placed at every $\frac{N}{k}$ layers in the $N$-layered LLM.
Image features obtained from self-supervised Vision Transformers (ViTs) [3, 41, 8] have been shown to exhibit rich semantic information at part-level granularity across similar, yet distinct object categories [1, 3, 71]. Motivated by these findings, we design a fusion module to extract such semantic information and facilitate part correspondence learning within the model. We define $\mathcal{E}$ to be the semantic extraction process using a self-supervised ViT [3] similarly to Amir et al. [1], and obtain semantic embeddings $\mathbf{X}_{\textnormal{semantic}}\in\mathbb{R}^{N_{I}\times S_{S}\times D_{S}}$, where $S_{S}$ and $D_{S}$ are the sequence length and hidden size of the semantic image embeddings, respectively. We then use $\mathbf{X}_{\textnormal{semantic}}$ as the key and value for a cross-attention extraction mechanism $\mathcal{A}$ with the queried EVA-CLIP embedding $\mathbf{X}_{\text{global}}$ to get semantic-rich global embeddings $\mathbf{X}^{\prime}_{\text{global}}$. This process is formalized by:
@@
Figure 4: Example image pairs in MixedParts with objects, common parts, and unique parts segmented and labeled. Each column represents a different image pair, derived from a set of diverse datasets with various levels of detail, PACO, PartImageNet, and ADE20K-Part-234, covering both rigid and non-rigid objects and parts. Each image pair is displayed across 3 rows to illustrate (i) the (possibly common or different) object(s), (ii) the common object part(s), and (iii) the unique object part(s) in each pair.
@@ 3.4 Correspondence Adaptation Module (CAM)
Due to the high cost of training LLMs with billions of parameters, many works have taken advantage of adaptive modules, with sizes merely a fraction of the LLMs’ original sizes, by freezing the LLMs and only training the modules for downstream tasks [17, 70, 18, 28]. These works have demonstrated strong performance and high efficiency while also avoiding the problem of catastrophic forgetting [49, 59]. In particular, the VPG-C module proposed by Li et al. [28] effectively exhibits the capability to adapt an LLM to multi-image reasoning while only accounting for $0.09\%$ of the entire model’s parameter count. Inspired by this finding, we leverage VPG-C for multi-image correspondence adaptation. Specifically, at select layers $l\in L$, we linearly project the last input token $\mathbf{t}^{l}_{S_{T}}\in\mathbb{R}^{D}$ via $f_{\textnormal{adaptation}}:\mathbb{R}^{D}\to\mathbb{R}^{D_{I}}$ to get an instruction-specific guidance embedding. We then use this embedding to enrich the query tokens $\mathbf{q}$ and obtain new query embeddings by:
@@
These context-guided tokens are finally used to query semantic- and context-rich visual information from $\mathbf{X}^{\prime}_{\text{global}}$ from the Correspondence Extraction Module to get:
@@ 3.5 Training Objective
Following past works [46, 24], the training loss $\mathcal{L}$ is composed of the next-token prediction loss $\mathcal{L}_{\textnormal{text}}$ and the segmentation mask loss $\mathcal{L}_{\textnormal{mask}}$. Here, $\mathcal{L}_{\textnormal{text}}$ is a causal cross-entropy (CE) loss computed from the predicted and right-shifted ground truth tokens $\mathbf{\hat{T}}^{K}$ and $\mathbf{y}$, while $\mathcal{L}_{\textnormal{mask}}$ combines a focal loss [31] and a DICE loss [40] derived from the predicted and ground truth masks $\mathbf{\hat{M}}$ and $\mathbf{M}$. The training objective can be summarized as follows:
$\mathcal{L}=\lambda_{\textnormal{text}}\mathcal{L}_{\textnormal{text}}+\mathcal{L}_{\textnormal{mask}}$
$\mathcal{L}_{\textnormal{test}}=\textnormal{CE}\left(\mathbf{\hat{T}}^{K},\mathbf{y}\right)$
$\mathcal{L}_{\textnormal{mask}}=\lambda_{\textnormal{focal}}\mathcal{L}_{\textnormal{focal}}\left(\mathbf{\hat{M}},\mathbf{M}\right)+\lambda_{\textnormal{Dice}}\mathcal{L}_{\textnormal{Dice}}\left(\mathbf{\hat{M}},\mathbf{M}\right),$
where $\lambda_{\textnormal{text}}$, $\lambda_{\textnormal{focal}}$, and $\lambda_{\textnormal{Dice}}$ represent the respective weighting coefficients for the loss components. These objectives enable an end-to-end training process focusing on optimizing the quality of both text and mask generation outputs.
@@ 4 MixedParts Dataset
Although multi-image datasets of various scales are available, they exhibit combinations of limitations, making them unsuitable for the part-focused semantic co-segmentation task. Limitations include the absence of fine-grained masks for segmentation [19, 28, 54, 55, 37], datasets being too small or domain-specific to facilitate generalizable LVLM training despite containing localized labels [2, 12, 52, 51, 62], or the lack of part-level information altogether [43, 51]. To address these challenges and enable effective training and evaluation of our part-focused semantic co-segmentation model, we introduce MixedParts, a novel dataset curated from publicly available datasets. Figure 4 provides examples from our dataset, demonstrating its diversity in object categories and visual details. We provide an overview of the public datasets utilized as part of MixedParts and detail the dataset construction process in Appendix A. Dataset statistics can be found in Appendix B.
@@ 5 Experiments
We evaluate Calico’s performance on the challenging MixedParts dataset, reporting the mean Intersection-over-Union (mIoU), Average Precision (AP50), and Recall, which assess the model’s segmentation performance. To evaluate its semantic label generation capability, following existing works [67, 7], we employ Semantic Similarity (SS) and Semantic IoU (S-IoU). Implementation details and descriptions of the evaluation metrics can be found in Appendix C. We perform evaluation on $\sim$1K image pairs, ensuring an equal distribution of image pairs from each original dataset and maintaining equal representation across all tasks. We report performance on all three subtasks of part-focused semantic co-segmentation – common objects, common parts, unique parts – and their average, which reflects overall performance on the MixedParts test set.
Baselines. To the best of our knowledge, we represent the first effort to tackle multi-image part-focused co-segmentation with part label identification. There is thus a lack of baselines for this new task, which we rectify by designing our own baselines from strong pretrained models in the semantic segmentation literature. Our baselines include a traditional zero-shot Mask R-CNN-based method and two finetuned LVLM-based approaches: (1) Multi-Image VLPart: VLPart [56] is an open-vocabulary part segmentation model which can identify object parts at different granularities. It utilizes a conventional Mask R-CNN [16] with a modern Swin Transformer backbone [34] and a CLIP [44] text classifier for open-world image classification. Although VLPart cannot perform co-segmentation, due to its strong zero-shot object-part segmentation capabilities, we perform segmentation on individual images and simply examine the common and unique predictions. Our implementation of Multi-Image VLPart, which ensures fairness, is detailed in Appendix C. (2) Multi-Image GLaMM: GLaMM [46] is a strong single-image segmentation-based LVLM constructed for the Grounded Conversation Generation (GCG) task for multi-round pixel-grounded conversations. The architecture incorporates a Vicuna-based backbone with SAM and a novel RoIAlign-based region encoder. Since GLaMM was not trained for multi-image processing, we replicate Calico’s multi-image implementation on the GLaMM codebase and finetune the mask decoder on MixedParts for GLaMM to learn multi-image reasoning, alongside LoRA, using the exact same hyperparameters. We initialize both Calico and Multi-Image GLaMM on GLaMM’s full model weights. [^1] https://huggingface.co/MBZUAI/GLaMM-FullScope. (3) Multi-Image LISA: LISA [24] is an LVLM trained to perform referring segmentation, built on a Vicuna-7B backbone and SAM, similarly to GLaMM. Likewise, we implement multi-image processing for LISA by sequentially feeding the mask decoder with the encoded images and corresponding segmentation tokens to obtain segmentation masks. We initialize LISA on the most recent 7B checkpoint [^2] https://huggingface.co/xinlai/LISA-7B-v1-explanatory. and finetune the mask decoder alongside LoRA.
Method mIoU AP50 Recall SS S-IoU
Multi-Image VLPart [56] 12.4 42.0 34.0 58.6 45.4
Multi-Image GLaMM [46] 30.3 51.6 40.8 63.9 56.3
Multi-Image LISA [24] 30.7 52.1 42.3 66.3 59.9
Calico (ours) 37.1 57.2 50.5 76.5 70.2
@@ 5.1 Experimental Results
Table 1 presents results comparing Calico against our baselines, Multi-Image VLPart, Multi-Image GLaMM, and Multi-Image LISA. Calico demonstrates superior performance on all metrics compared to all zero-shot and finetuned approaches, achieving relative gains of 20.8%, 9.8%, and 19.4% on segmentation-based metrics mIoU, AP50, and Recall, respectively, and 15.4% and 17.2% on text-based metrics SS and S-IoU. Although VLPart demonstrates strong zero-shot single-image object-part segmentation performance, it struggles on our multi-image tasks, even on text-based metrics where the model has direct access to the ground truth class labels. When finetuned on MixedParts, Multi-Image GLaMM and LISA exhibit performance improvements compared to the zero-shot VLPart baseline. However, GLaMM still significantly lags behind Calico’s performance despite our model being initialized from the same weights, demonstrating the effectiveness of our proposed modules, which we further ablate in Section 5.2. We present qualitative examples in Figure 2 and additional experiments in Appendix E.
(a) Common Object (b) Common Parts (c) Unique Parts
The common object is the person. The images show a snake and a dog. The images contain a bed and a table.
The object present in the images is the car. The common parts in both objects are the body and the head. The unique parts between the objects are a headboard, a top, and a leg.
Table 2: Calico Qualitative Results. Across various object types, Calico can segment (a) common objects in the images, (b) shared parts between objects from different classes, and (c) distinct parts unique to each object.
@@ 5.2 Ablations
Calico Components. To validate the effectiveness of the individual components of CALICO, we conduct an ablation that isolates the impact of the Correspondence Extraction Module (CEM) and the Correspondence Adaptation Module (CAM). In Figure 5, we report results of Calico without the CEM module (w/o CEM), without the CAM module (w/o CAM), or removing both (w/o CEM w/o CAM). Specifically, the Calico variant without the CEM module utilizes only the image embeddings for fusion with the LVLM outputs’ last hidden states, while the Calico variant without the CAM module only injects DINOv2 features without fusing them with the last hidden states.
w/o CEMw/o CAMw/o CEM w/o CAMCALICO020406080 $28.20$ $33.40$ $31.30$ $37.10$ $50.00$ $54.80$ $53.00$ $57.20$ $35.50$ $47.20$ $42.50$ $50.50$ $55.40$ $74.50$ $68.10$ $76.50$ $46.80$ $67.90$ $61.10$ $70.20$ Metric Scores (%)mIoUAP50RecallSSS-IoU
Figure 5: Ablations on the Correspondence Extraction and Correspondence Adaptation Modules.
@@
Ablation results indicate that the proposed modules play a significant role in Calico’s ability to accurately co-segment and identify common objects, common parts, and unique parts. Specifically, removing CEM results in a notable decrease across all metrics. Similarly, excluding the CAM module leads to a degradation w.r.t. segmentation and labeling performance. CEM substantially contributes to the segmentation performance as this module enables the model to learn semantic relationships in multi-image contexts. Interestingly, keeping CAM while removing CEM results in degraded performance compared to the version without both modules. This may mean that CAM does not work well when there is no external signal from CEM, in which case it mostly duplicates the image features in the input and potentially confuses the model. However, when CEM is included, the variant without both modules underperforms the one without CAM, which in turn is outperformed by Calico. This demonstrates that CAM is beneficial when there are external semantic signals for the part-focused semantic co-segmentation task.
@@
Calico Injecting CAMs. We perform further ablations to examine the efficacy of injecting 2 evenly spaced Correspondence Adaptation Modules (CAM) into our Calico LVLM, and present results in Table 3. Past works [64, 28] have demonstrated the effectiveness of using or injecting information at the intermediate layer $\frac{K}{2}$ as guidance for learning. Since our task involves multimodal understanding at multiple granularities (object and part), we use 2 evenly spaced layers to incorporate semantic features at different levels of the LVLM’s learning, encouraging the model to focus on various object-part correspondences. This results in the best segmentation performance, as shown in Table 3, compared to the 1-layer or 3-layer variants. Although injecting 3 adaptive modules interestingly improves results on our text metrics, this comes at the expense of segmentation performance; therefore, we opt for injecting 2 layers to preserve grounding performance.
Layers mIoU AP50 Recall SS S-IoU
16 33.7 54.8 47.3 73.6 66.7
11, 22 (Calico) 37.1 57.2 50.5 76.5 70.2
8, 16, 24 36.4 56.7 50.2 77.4 70.8
Table 3: Ablations on CAM layer injection.

### v2

@@ 3.3 Correspondence Extraction Module (CEM)
Figure 4: Overview of our Correspondence Extraction and Correspondence Adaptation Modules. In Calico, $k$ CAMs are placed at every $\frac{N}{k}$ layers in the $N$-layered LLM.
Image features obtained from self-supervised Vision Transformers (ViTs) [5, 45, 11] have been shown to exhibit rich semantic information at part-level granularity across similar, yet distinct object categories [2, 5, 75]. Motivated by these findings, we design a fusion module to extract such semantic information and facilitate part correspondence learning within the model. We define $\mathcal{E}$ to be the semantic extraction process using a self-supervised ViT [5] similarly to [2], and obtain semantic embeddings $\mathbf{X}_{\textnormal{semantic}}\in\mathbb{R}^{N_{I}\times S_{S}\times D_{S}}$, where $S_{S}$ and $D_{S}$ are the sequence length and hidden size of the semantic image embeddings, respectively. We then use $\mathbf{X}_{\textnormal{semantic}}$ as the key and value for a cross-attention extraction mechanism $\mathcal{A}$ with the queried EVA-CLIP embedding $\mathbf{X}_{\text{global}}$ to get semantic-rich global embeddings $\mathbf{X}^{\prime}_{\text{global}}$. This process is formalized by:
@@
Figure 5: Example Image Pairs in MixedParts with Common Objects, Common Parts, and Unique Parts segmented and labeled. Each column represents a different image pair, derived from a set of diverse datasets with various levels of detail, PACO-LVIS, PartImageNet, and ADE20K-Part-234, covering both rigid and non-rigid objects and parts. Each image pair is displayed across 3 rows to illustrate (i) the (possibly common or different) object(s), (ii) the common object part(s), and (iii) the unique object part(s) in each pair.
@@ 3.4 Correspondence Adaptation Module (CAM)
Due to the high cost of training LLMs with billions of parameters, many works have taken advantage of adaptive modules, with sizes merely a fraction of the LLMs’ original sizes, by freezing the LLMs and only training the modules for downstream tasks [20, 74, 21, 31]. These works have demonstrated strong performance and high efficiency while also avoiding catastrophic forgetting [53, 63]. In particular, VPG-C [31] effectively exhibits the capability to adapt an LLM to multi-image reasoning while only accounting for $0.09\%$ of the entire model’s parameter count. Inspired by this finding, we leverage VPG-C for multi-image correspondence adaptation. Specifically, at select layers $l\in L$, we linearly project the last input token $\mathbf{t}^{l}_{S_{T}}\in\mathbb{R}^{D}$ via $f_{\textnormal{adaptation}}\!:\!\mathbb{R}^{D}\to\mathbb{R}^{D_{I}}$ to get an instruction-specific guidance embedding. We then use this embedding to enrich the query tokens $\mathbf{q}$ and obtain new query tokens by:
@@
These context-guided query tokens are finally used to extract semantic- and context-rich visual information from $\mathbf{X}^{\prime}_{\text{global}}$ from the Correspondence Extraction Module via:
@@ 3.5 Training Objective
Following prior work [50, 27], we optimize a combined training loss $\mathcal{L}\!=\!\lambda_{\textnormal{text}}\mathcal{L}_{\textnormal{text}}+\mathcal{L}_{\textnormal{mask}}$, where $\mathcal{L}_{\text{text}}$ is the next-token prediction loss and $\mathcal{L}_{\text{mask}}$ is the segmentation loss. The weights $\lambda_{\text{text}}$, $\lambda_{\text{focal}}$, and $\lambda_{\text{Dice}}$ control the contribution of each loss component. Here, $\mathcal{L}_{\textnormal{text}}$ is a causal cross-entropy (CE) loss computed from the predicted and right-shifted ground truth tokens $\mathbf{\hat{T}}^{K}$ and $\mathbf{y}$,
$\mathcal{L}_{\textnormal{text}}=\textnormal{CE}\left(\mathbf{\hat{T}}^{K},\mathbf{y}\right),$
while $\mathcal{L}_{\textnormal{mask}}$ combines a focal loss [35] and a DICE loss [44] derived from predicted and ground truth masks $\mathbf{\hat{M}}$ and $\mathbf{M}$,
$\mathcal{L}_{\textnormal{mask}}=\lambda_{\textnormal{focal}}\mathcal{L}_{\textnormal{focal}}\left(\mathbf{\hat{M}},\mathbf{M}\right)+\lambda_{\textnormal{Dice}}\mathcal{L}_{\textnormal{Dice}}\left(\mathbf{\hat{M}},\mathbf{M}\right).$
@@ 4 MixedParts Dataset
Although multi-image datasets of various scales are available, they present combinations of limitations that make them unsuitable for the part-focused semantic co-segmentation task. These include the absence of fine-grained masks for segmentation [22, 31, 58, 59, 41], datasets being too small or domain-specific to facilitate generalizable LVLM training despite containing localized labels [3, 15, 56, 55, 66], or the lack of part-level information altogether [47, 55]. To address these gaps, we introduce MixedParts, a new dataset curated from publicly available sources to support training and evaluation of part-focused semantic co-segmentation models. Figure 5 provides examples from our dataset, highlighting its diversity in object categories and visual details. Appendix A describes the dataset construction process and source datasets, while Appendix B summarizes dataset statistics.
@@ 5 Experiments
(a) Common Object (b) Common Parts (c) Unique Parts
The common object is the person. The images show a snake and a dog. The images show a bed and a pool table.
The images include a car. The detected common parts are a body and a head. The unique parts present are a headboard, a bed, and a pocket.
Figure 6: Qualitative Results. Calico can co-segment (a) common objects, (b) shared parts between objects from different classes, and (c) distinct parts unique to each object. Our model demonstrates pixel-grounded understanding of various non-rigid (e.g., humans, animals) and rigid objects (e.g., car, bed), including less common objects and their parts (e.g., the bed and pocket of a pool table).
Metrics. We evaluate Calico’s performance on the MixedParts dataset, reporting the Average Precision (AP50), mean Intersection over Union (mIoU), and Recall, which assess the model’s segmentation performance. To evaluate its semantic label generation capability, following existing works [71, 9], we employ Semantic Similarity (SS) and Semantic IoU (S-IoU). Implementation details and descriptions of the evaluation metrics can be found in Appendix C.2. We perform evaluation on $\sim$1K image pairs, ensuring an equal distribution of image pairs from each original dataset and maintaining equal representation across all tasks. We report performance on all three subtasks of part-focused semantic co-segmentation – common objects, common parts, unique parts – and their average, which reflects overall performance on the MixedParts test set.
Baselines. To the best of our knowledge, this is the first effort to tackle multi-image part-focused co-segmentation with part label identification. There is thus a lack of baselines for this new task, which we rectify by designing our own baselines from strong pretrained models in the semantic segmentation literature. Our baselines comprise three zero-shot approaches: one modular method (Cascade) that combines pretrained models [22, 1, 27], each excelling at a different aspect of our novel task, and two methods that leverage state-of-the-art part segmentation models to compare their outputs (PartGLEE [32] and VLPart [60]). Additionally, we report results for two finetuned segmentation-based LVLMs (GLaMM [50] and LISA [27]). Implementation details for all models are available in Appendix C.
Method AP50 mIoU Recall SS S-IoU
Cascade [22, 1, 27] 5.7 27.9 19.0 32.2 14.8
Multi-Image PartGLEE [32] 1.2 29.3 9.7 78.5 63.3
Multi-Image VLPart [60] 13.4 42.8 34.6 59.1 46.5
Multi-Image GLaMM [50] 42.9 59.9 54.9 76.8 71.2
Multi-Image LISA [27] 41.4 59.7 55.5 78.7 72.5
Calico 45.9 63.7 59.7 82.7 77.1
@@ 5.1 Experimental Results
Table 1 presents results comparing Calico against our baselines. Calico demonstrates superior performance on all metrics compared to all zero-shot and finetuned approaches, achieving relative gains of 7.0%, 6.3%, and 7.6% on segmentation-based metrics AP50, mIoU, and Recall, respectively, and 5.0% and 6.3% on text-based metrics SS and S-IoU. These performance gains come with no overhead; in fact, as shown in Figure 2, Calico demonstrates high efficiency compared to LISA and GLaMM in terms of TFLOPS and inference time. Details can be found in Appendix D.1.
The cascade method fails to perform well on our task, highlighting the task’s challenging nature and suggesting that it cannot be addressed merely by combining standalone modules. Although PartGLEE and VLPart demonstrate strong zero-shot single-image object-part segmentation performance, they struggle with our multi-image tasks. In particular, PartGLEE produces 29 object predictions per image on average, increasing its likelihood of matching ground truth labels; yet Calico’s text scores are superior despite focusing on a single most relevant object in context. When finetuned on MixedParts, Multi-Image GLaMM and LISA exhibit performance improvements compared to the zero-shot baselines. However, GLaMM still lags behind Calico’s performance despite both initialized from the same weights, demonstrating the effectiveness of our proposed modules, which we further ablate in Section 5.2.
Qualitative results in Figure 6 showcase Calico’s strong pixel-level understanding across all three tasks in part-focused semantic co-segmentation. Our model effectively localizes both non-rigid objects (e.g., humans, animals) and rigid ones (e.g., cars, beds), and demonstrates part-level reasoning even for less common object classes and their parts, such as the bed and pocket of a pool table. Furthermore, Figure 7 illustrates Calico’s contextual understanding, where different image pairings prompt the model to segment different objects accordingly, rather than defaulting to the most salient object in each image. Additional experiments are detailed in Appendix D.
[Uncaptioned image] [Uncaptioned image]
The images show a dog.
[Uncaptioned image] [Uncaptioned image]
The images include a car.
Figure 7: CALICO In Context. Calico outputs are highly context-driven when distinguishing objects across images, despite variations in angle, size, saliency, etc.
@@ 5.2 Ablations
Calico Components. We perform ablations to assess the contribution of each component in Calico, including the Correspondence Extraction Module (CEM) and the Correspondence Adaptation Module (CAM), alongside Q-Former and DINOv2. Table 2 reports results for Calico variants: without Q-Former (w/o Q-Former), without DINOv2 in CEM (w/o DINO), without CEM (w/o CEM), without CAMs (w/o CAM), or removing both (w/o CEM w/o CAM). Specifically, the w/o CEM variant fuses each LVLM layer’s image embeddings with only the previous layer’s last hidden state, while w/o CAM injects CEM features without fusing them with the last hidden state. In the w/o DINO setting, no cross-attention is applied, as there are no DINOv2 features to fuse with EVA-CLIP. Since our CEM and CAM modules are designed for models with Q-Formers, we approximate the w/o Q-Former variant using a linear projection, similar to the linear ablation in [31].
@@
Ablation results show that each proposed module contributes to Calico’s ability to accurately co-segment and identify common objects, common parts, and unique parts. Firstly, removing Q-Former – and thus disabling most of CAM and CEM – leads to a substantial drop in performance compared to Calico, underscoring its importance within our model architecture. Removing CEM results in a decrease in performance across all metrics relative to Calico and demonstrates comparable performance to the “w/o DINO” variant, highlighting the critical role of both DINOv2 and CEM in Calico. CEM substantially contributes to segmentation performance as this module enables the model to learn semantic relationships in multi-image contexts. Excluding the CAM modules leads to a decline in labeling performance, suggesting that CAM facilitates the integration of image-derived features into the language space. Interestingly, keeping CAM while removing CEM results in worse performance than removing both, indicating that CAM may introduce redundant or uninformative features when not guided by external signals from CEM. In contrast, when CEM is present, removing CAM still outperforms the variant without both modules, and full Calico achieves the best results. This demonstrates that CAM is most beneficial when there are external semantic signals for the part-focused semantic co-segmentation task.
@@
Calico Injecting CAMs. We perform further ablations to examine the efficacy of injecting 2 evenly spaced Correspondence Adaptation Modules (CAM) into our Calico LVLM, and present results in Figure 8. Past works [68, 31] have demonstrated the effectiveness of using or injecting information at the intermediate layer $\frac{K}{2}$ as guidance for learning. Since our task involves multimodal understanding at multiple granularities (object and part), we use 2 evenly spaced layers to incorporate semantic features at different levels of the LVLM’s learning, encouraging the model to focus on various object-part correspondences. This results in the best segmentation and labeling performance compared to the 1-layer or 3-layer variants.
1611+22 (CALICO)8+16+24020406080 $44.7$ $45.9$ $43.8$ $63.2$ $63.7$ $61.4$ $58.6$ $59.7$ $57.2$ $82.3$ $82.7$ $79.7$ $76.8$ $77.1$ $74.3$ Metric Scores (%)AP50mIoURecallSSS-IoU
Figure 8: Ablations on CAM layer injection.
Setting AP50 mIoU Recall SS S-IoU
w/o Q-Former 38.5 59.2 44.8 64.5 55.6
w/o DINO 43.9 61.7 57.1 80.2 74.5
w/o CEM 43.6 61.6 57.5 80.8 75.2
w/o CAM 45.9 63.3 59.7 82.0 76.5
w/o CEM w/o CAM 44.1 62.7 58.1 81.6 76.3
Calico 45.9 63.7 59.7 82.7 77.1
Table 2: Ablations on Calico Components, including Q-Former, DINOv2, the Correspondence Extraction Module, and the Correspondence Adaptation Modules.